OpenStudy (geometry_hater):

Solve for x 3X^2 - 2x + 5 = 0

OpenStudy (anonymous):

Which method do you have to use??? Factorisation??

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (geometry_hater):

ok it just says solve the equation graphically

OpenStudy (anonymous):

Put different values of x and get corresponding values of y, plot them and see where the graph intersects the x-axis....... the intersection points are the solutions....

OpenStudy (anonymous):